OpenStudy (karolynxx):

A system of equations has no solution. If y = 8x + 7 is one of the equations, which could be the other equation? y = 8x + 7 y = 8x – 7 y = –8x + 7 y = –8x – 7

OpenStudy (welshfella):

Hint:- Equations that form lines parallel to each other have no solution Parallel means they have the same slope. consider the general form y = mx + b - m is the value of the slope.

OpenStudy (welshfella):

which equation has the same slope as y = 8x + 7?

OpenStudy (welshfella):

ignore y = 8x + 7 as it is the same equation as in the question

OpenStudy (welshfella):

compare y = 8x + 7 with the other 3 choices

OpenStudy (karolynxx):

i still dont understand

OpenStudy (mayankdevnani):

is it multiple choice question ?

OpenStudy (karolynxx):

yes the choices are at the bottom

OpenStudy (mayankdevnani):

i mean there are more than one correct answers

OpenStudy (karolynxx):

the anwsers are a.y = 8x + 7 b.y = 8x – 7 c.y = –8x + 7 d.y = –8x – 7

OpenStudy (karolynxx):

cant be more then one

OpenStudy (mayankdevnani):

but it should be more than one

OpenStudy (mayankdevnani):

so i'll tell you method how to solve this type of question

OpenStudy (welshfella):

@karolynxx - just check out the number which multiplies x in the case of y = 8x + 7 its 8 what other equation has 8x in it?

OpenStudy (welshfella):

- its easier than you think